OpenStudy (anonymous):

the limiting sum of a geometric series is 5 and the 2nd term is 6/5. find the first term and the common ratio.

OpenStudy (rane):

sorry bt can u say again wht's the role of 5

OpenStudy (anonymous):

its the limiting sum

OpenStudy (anonymous):

 so its that

OpenStudy (anonymous):

thats the limiting sum formula

OpenStudy (rane):

i'm doing geometric series bt haven't yet came across limiting sum!

OpenStudy (anonymous):

you should come across it soon

OpenStudy (rane):

ahan

OpenStudy (jhannybean):

Now if you're given the second term,which is 6/5, hw would you use this to find the first term in the sequence? :)

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

what the name of this formula ?

OpenStudy (jhannybean):

geometric series.

hartnn (hartnn):

like first term is \(a_1\), 2nd term, you just multiply by r, 2nd term = \(a_1r\), 3rd term, you again multiply by r, 3rd term = \(a_1r^2\) terms ----> \(a_1,a_1r,a_1r^2,a_1r^3....\) so, general term = \(a_n=a_1r^{n-1}\) here you have ar = 6/5 and a/(1-r) =5 solv simultaneously...

OpenStudy (jhannybean):

hmm...

hartnn (hartnn):

a= 6/(5r)

OpenStudy (jhannybean):

ar = 6/5 therefore a = 6/5r lol..... -_- \[\huge \frac{\frac{6}{5r}}{1-r}=5\]

hartnn (hartnn):

yes^

hartnn (hartnn):

@yasmin95 could u solve now ?

OpenStudy (anonymous):

yes thanks :)

hartnn (hartnn):

welcome ^_^

OpenStudy (jhannybean):

\[\large \frac{6}{5r}= 5-5r\]\[\large 5r + \frac{6}{5r}=5\] Im too l azy.

hartnn (hartnn):

quadratic equation...i guess yasmin already solved it :)

OpenStudy (jhannybean):

I don't even know.....how to.... :| havent bothered, lol.

OpenStudy (jhannybean):

@yasmin95, how did you solve thiss